Suppose the random variable X has the probability density function (p.d.f) given as
f(x) = c(1 − x^2), for − 1 < x < 1
0, elsewhere
for some constant c.
(i) What is the value of c?
(ii) Derive the cumulative distribution function of X, F(X).
(iii) Hence or otherwise, find P(X > 0.7).
1
Expert's answer
2020-04-02T13:49:12-0400

"f(x)=\begin{cases}
c(1-x^2), \ \ -1<x<1
\\
0, \ \ \text{ elsewhere }
\end{cases}"

(i) To find the value of "c" , we can use the fact that "\int \limits _{-\infin}^{\infin}f(x)dx=1"

"\int \limits _{-\infin}^{\infin}f(x)dx= \int \limits _{-1}^{1}c(1-x^2)dx =c(x-\frac{x^3}{3}) \bigg| _{x=-1}^{x=1}=
\frac{4c}{3}=1 \ \Rightarrow c=3/4"


Answer: "c=3\/4"


(ii) "F(x)=\int \limits _{-\infin}^{x}f(t)dt"

if "x\leq -1:" "F(x)= \int \limits _{-\infin}^{x}f(t)dt= \int \limits _{-\infin}^{x}0dt=0"

if "-1<x<1: F(x)= \int \limits _{-\infin}^{x}f(t)dt= \int \limits _{-1}^{x}\frac{3}{4}(1-t^2)dt=\frac{3}{4}(t-\frac{t^3}{3})-\frac{3}{4}(-1-\frac{(-1)^3}{3})="

"= \frac{3}{4}t-\frac{1}{4}t^3+\frac{1}{2}"

if "x\geq 1: F(x)=\int \limits _{-\infin}^{x}f(t)dt= \int \limits _{-1}^{1}f(t)dt+ \int \limits _{1}^{x}f(t)dt=\ 1+0=1"


Answer: "F(x)=\begin{cases}
0, \ \ x\leq -1
\\
\frac{3}{4}t-\frac{1}{4}t^3+\frac{1}{2} , \ \ -1<x<1
\\
1, \ \ 1\leq x
\end{cases}"

(iii) "P(X>0.7)=1-P(X\leq 0.7)=1-F(0.7)=1-(\frac{3}{4}\times 0.7-\frac{1}{4}\times 0.7^3+\frac{1}2)="

"=0.06075"


Answer: "P(X>0.7)=0.06075"